Ella has only nickels (x) and dimes (y) in her piggy bank. She has 36 coins in all and has four more dimes than nickels. Find how much money she has in her piggy bank using a 2 variable, 2 equation system solution.
Answer by ikleyn(52817)

.

From the condition, you have two equations


    y + x = 36    (1)    (total coins)

    y - x =  4    (2)    (four more dimes than nickels)


To solve the system, add the equations (both sides). The terms "x" and "-x" will cancel each other,

leaving you one equation for unknown y

    
    2y = 36 + 4 = 40,


from which you get


     y = 40/2 = 20.


So, there are 20 dimes.


Now substitute y= 20 into equation (1). You will get


    20 + x = 36,


hence


    x = 36 - 20 = 16.


ANSWER.  x = 16 (nickels)  and y = 20 (dimes).

Solved and thoroughly explained.